# Multi S-Graphs: an Efficient Real-time Distributed Semantic-Relational Collaborative SLAM

### 10 Jan 2024 v1

@@ Abstract
Collaborative Simultaneous Localization and Mapping (CSLAM) is critical to enable multiple robots to operate in complex environments. Most CSLAM techniques rely on raw sensor measurement or low-level features such as keyframe descriptors, which can lead to wrong loop closures due to the lack of deep understanding of the environment. Moreover, the exchange of these measurements and low-level features among the robots requires the transmission of a significant amount of data, which limits the scalability of the system. To overcome these limitations, we present Multi S-Graphs, a decentralized CSLAM system that utilizes high-level semantic-relational information embedded in the four-layered hierarchical and optimizable situational graphs for cooperative map generation and localization while minimizing the information exchanged between the robots. To support this, we present a novel room-based descriptor which, along with its connected walls, is used to perform inter-robot loop closures, addressing the challenges of multi-robot kidnapped problem initialization. Multiple experiments in simulated and real environments validate the improvement in accuracy and robustness of the proposed approach while reducing the amount of data exchanged between robots compared to other state-of-the-art approaches.
Software available within a docker image: https://github.com/snt-arg/multi_s_graphs_docker
@@ I Introduction
Collaborative Simultaneous Localization and Mapping (CSLAM) is a technology that enables multi-robot systems to work together to create a map of their environment while simultaneously determining their own positions within the map. Most CSLAM techniques, such as [1], [2], [3], are highly based on the transmission of low-level features, such as keyframe descriptors, between robots. With these low-level features, each robot creates its own map and can match it to integrate them into a common map. This is used for both visual and LiDAR-based approaches. However, the use of these low-level features usually leads to incorrect loop closures [4]. Works such as [1] or [5] focus on robustifying their loop closure algorithms to avoid incorrect loop closures. The main drawback of these methods emerges from the fact that the system has limited understanding of the environment.
Novel approaches to single-robot SLAM address this issue of limited understanding of the environment by incorporating meaningful metric-semantic-relational information, with works like Hydra [6] which exploits 3D scene graphs for loop closure, or S-Graphs+ [7], [8], which couple in a situational graph modeled as a single optimizable factor graph, the 3D scene graph with the underlying SLAM graph. These works demonstrate that extracting semantic meaning from metric information and incorporating it with appropriate relational constraints such as walls, rooms, floors, etc., improve the overall precision of underlying SLAM while robustifying loop closures and generating meaningful environmental maps. Recently, Hydra-Multi [9] and CURB-SG [10] took advantage of 3D scene graphs for collaborative SLAM in a centralized way. These innovative approaches harness semantic-relational knowledge to exchange higher-level information between the agents to fortify inter-robot loop closures. However, these methods, apart from being centralized, explicitly add loop closure constraints to underlying robot poses instead of implicitly constraining the same semantic entities detected by different agents.
Thus in this work, we present Multi S-Graphs, a LiDAR-based distributed CSLAM system harnessing the potential of the multi-layered Situational Graphs (S-Graphs), a single optimizable factor graph that combines a pose graph (i.e. with keyframes) and a 3D scene graph with semantic-relational concepts, to generate precise environmental maps and locating the different robots in this map while exchanging minimum information between robots. To the best of our knowledge, we present the first real-time distributed collaborative SLAM system based on optimizable situational graphs. Fig. 1 shows the Multi S-Graphs executed on a real experiment.
@@
    A hybrid room-based descriptor combining fine-grained point cloud information with semantic knowledge to identify inter-robot loop closures.
@@ II-A Multi-Agent SLAM
There are two main categories of multi-agent SLAM architectures: centralized and decentralized.
@@
Figure 2: Multi S-Graph architecture viewed from $Robot_{i}$ perspective. Odometry measures and 3D Lidar data are the inputs that the robotic platform provides. Each robot executes its own Local S-Graph system and generates the Room Descriptors and the Distilled S-Graphs that are exchanged with the rest of the robotic network. Messages exchanged between robots are represented in gray. From the point of view of the $Robot_{i}$ the rest of the agents provide the Distilled graphs and the Room descriptors that are used to generate the Collaborative S-Graph. The information provided by the rest of the robots appears in green color.
@@
II-B Multi-Agent 3D Scene Graphs
@@ III System Overview
1. Local S-Graph Generation (Section IV). To generate the local S-Graph for each robot, we use [8] which runs onboard each one using the odometry and 3D LiDAR measurements of the robot. It also estimates the position and trajectory followed by the robot while generating it.
@@
3. Room-Based Inter-Robot Loop Closure (Section VI). Relying only on the generated semantic graph for loop closure can generate alignment errors in the presence of symmetry in the environment. In order to combine the fine-grained information of the point clouds and take advantage of the high-level semantic information that each room contains, we generate a hybrid descriptor for each room, a Room Descriptor. This module is in charge of two tasks: generating the descriptors and finding appropriate room matches.
@@ VI-A Room descriptor generator
To generate these descriptors, we rely on the Scan Context descriptor [27]. This is an egocentric, yaw-invariant, point cloud descriptor that has demonstrated state-of-the-art results, given its simplicity and fast convergence time. The main drawback of this descriptor is that it is sensitive to linear motions. To solve this, we take advantage of the semantic information in the room by generating a Scan Context descriptor from the center of each room, avoiding translation errors. To generate the room descriptor, we need a Room Centric point cloud, which is a point cloud built by combining all the point clouds of the keyframes within a room. Each room keyframe $R_{k_{i}}$ can be expressed as:
@@ VI-B Loop Closure detection
Fine Alignment: Whenever a match is found between the robot and other agents’ room, it tries to obtain an improved transform using the Room Centric Point cloud that originates the Scan Context. For this, the robot asks for the down-sampled point cloud and uses a VGICP[28] registration algorithm. The validity of the relative transform is determined by the alignment distance and the matching threshold $t_{d}$.
If a match is found, the room match with the relative transformation between rooms is sent into the following module (Section V) to connect the different graphs.
@@ VII Collaborative Situational Graph
Collaborative Graph: If a room match is found using the room matching step (Section VI), we formulate the factors between the local s-graph and the distilled s-graphs in the 2 steps. In the first step, a room-to-room factor is created for the matched rooms. For the given matched rooms, lower-level wall-to-wall factors are established, which robustify the alignment of the graphs. Whenever local S-Graphs and external distilled S-Graphs are connected with appropriate factors, a common optimization is performed similarly to [8]. After this optimization, the external graphs are aligned with the robot’s local coordinate frame so that the information provided by the other robots is integrated into the collaborative S-graph, using the same coordinated frame as the local robot for representing the different entities of the external distilled graphs, as can be seen in Fig. 1.
@@ VIII Experimental Validation
VIII-A Methodology
(a) C1F0s
(b) C1F2s
(c) SE1
Figure 4: Simulated experiments, showing the trajectory and the point cloud map generated for each robot
@@ VIII-A Methodology
Real-world Data. The datasets utilized in our experiments consist of a collection of in-house datasets acquired from a construction site: C1F1, C1F2, C3F1, C3F2, and C2F2. In addition, an experiment with 3 robots was performed on the C2F2 stage enabled by the size of the scenario. To ensure a fair comparison among the methods, we leverage the robot encoder odometry data, which is available due to the presence of physical encoders on the robot, and apply it consistently across all approaches. Unfortunately, our real-world datasets do not include IMU data and consequently, algorithms that rely heavily on IMU, such as DCL-SLAM and Disco-SLAM, could not be evaluated.
@@
(a) C1F1
(b) C1F2
(c) C3F1
(d) C3F2
(e) C2F2
VIII-B Experimental Setup
The robot used for generating the datasets is a Boston Dynamics Spot with a Velodyne VLP-16 onboard. The experiments were carried out using a laptop computer with an Intel i7-10870H (8 cores, 2.2 GHz) with 32 GB of RAM memory. Multi S-Graphs is developed under S-Graphs with some code refactoring in the implementation and uses g2o library [30] for the factor graphs implementation and ROS and ROS 2 nodes for implementing the algorithms.
VIII-C Results and Discussions
Simulated Data. Fig. 4 shows the qualitative results of Multi S-Graphs obtained during the simulated experiments. The outcomes presented in Table I show that Multi S-Graph algorithm exhibits a lower Average Trajectory Error (ATE) for the aggregated robot trajectory compared with the selected state-of-the-art (SOTA) CSLAM algorithms. Our approach outperformed all other decentralized algorithms, reducing the accuracy error by more than $90\%$ ($93.7\%$, $94.5\%$, and $98.4\%$ for DCL-SLAM, Disco-SLAM, and Swarm-SLAM respectively). It also achieved better accuracy than LAMP 2.0, with $72.9\%$ less error, which requires precise initial poses of the robots. The lack of features in the simulation environment causes measurement errors to adversely affect optimization. The extraction of semantic information specific to this type of environment by our algorithm overcomes this source of error.
@@ VIII-C Results and Discussions
Ablation Study. The results of this study are collected in Table V. We can see that the utilization of hierarchical optimization, which establishes relationships between the walls of two rooms among agents (Rooms and Planes), means a significant improvement of accuracy, reducing the ATE by $25.6\%$ with an increase in the amount of data by $77\%$. On the other hand, the use of Fine alignment improves the performance by $73.6\%$ and $25.6\%$ over the Rooms and Rooms and Walls configurations, respectively, with a 40 times higher data exchange requirement. In general, fine alignment techniques perform better compared to their non-fine alignment counterparts, primarily due to the limited angular resolution of the Scan Context, which is insufficient for precise alignment (see Section VI). Furthermore, the use of wall-to-wall factors contributes to alignment improvement, consequently significantly enhancing overall performance. Consequently, the best performance is obtained by combining the semantic information extracted from the walls with fine alignment using the Room Centric point clouds, Rooms and Walls with Fine alignment, improving the performance by up to $74.7\%$. However, due to the larger amount of data exchanged required, when communication channels are limited, Rooms and Walls, which only use semantic information, may be an interesting solution. This makes it possible to avoid sending heavy point cloud messages without a significant decrease in accuracy.
@@
Figure 6: Accumulated data shared between agents, consisting of Walls Nodes (Wall GN) and Room Nodes (Room GN) for the Semantic Graph, and the Room Scan Contexts (Room SC) and Room Centric point clouds (Room PC) that form the Room Descriptors.
@@ IX Conclusions and Future Work
In this work a distributed CSLAM system is presented, leveraging the semantic-relational features extracted by the S-Graphs [7], [8], in order to filter and reduce the amount of data that has to be transmitted between robots while outperforming state-of-the-art mapping and positioning accuracy in structured environments.
We tested our algorithm for a map generation task in multiple scenarios, simulated and real, outperforming other state-of-the-art lidar-based collaborative SLAM algorithms in terms of precision and bandwidth, reducing the trajectory and map reconstruction error by more than $90\%$ while saving $95\%$ of the data exchanged between robots. As our approach reduces the amount of data exchanged between agents, it makes it suitable for low-bandwidth scenarios, as well as for a large number of robots using the same communication channel.
The main drawback is the dependence on semantic-relational structures, defined in our case as four-walled rooms existing in building-like environments, which are not always easy to extract.
For future work, a more generic way of finding alignment between Situational Graphs should be found to increase the variety of scenarios in which Multi S-Graphs can be used. Including other sensory input for the generation of the S-Graph and the descriptors is required to allow the use of this system for heterogeneous robotic systems.

### 10 Apr 2024 v2

# Multi S-Graphs: An Efficient Distributed Semantic-Relational Collaborative SLAM

@@ Abstract
Collaborative Simultaneous Localization and Mapping (CSLAM) is critical to enable multiple robots to operate in complex environments. Most CSLAM techniques rely on raw sensor measurement or low-level features such as keyframe descriptors, which can lead to wrong loop closures due to the lack of deep understanding of the environment. Moreover, the exchange of these measurements and low-level features among the robots requires the transmission of a significant amount of data, which limits the scalability of the system. To overcome these limitations, we present Multi S-Graphs, a decentralized CSLAM system that utilizes high-level semantic-relational information embedded in the four-layered hierarchical and optimizable situational graphs for cooperative map generation and localization in structured environments while minimizing the information exchanged between the robots. To support this, we present a novel room-based descriptor which, along with its connected walls, is used to perform inter-robot loop closures, addressing the challenges of multi-robot kidnapped problem initialization. Multiple experiments in simulated and real environments validate the improvement in accuracy and robustness of the proposed approach while reducing the amount of data exchanged between robots compared to other state-of-the-art approaches.
Software as docker image: https://github.com/snt-arg/multi_s_graphs_docker
Video: https://youtu.be/0JsDn651rdc
@@ I Introduction
Collaborative Simultaneous Localization and Mapping (CSLAM) is a technology that enables multi-robot systems to work together to create a map of their environment while simultaneously determining their own positions within the map. Most CSLAM techniques, such as [1],[2], [3], are highly based on the transmission of low-level features, such as keyframe descriptors, between robots. With these low-level features, each robot creates its own map and can match it to integrate them into a common map. This is used for both visual and LiDAR-based approaches. However, the use of these low-level features usually leads to incorrect loop closures [4]. Works such as [1], [5] focus on robustifying their loop closure algorithms to avoid incorrect loop closures. The main drawback of these methods emerges from the fact that the system has limited understanding of the environment.
Novel approaches to single-robot SLAM address this issue of limited understanding of the environment by incorporating meaningful metric-semantic-relational information, with works like Hydra [6] which exploits 3D scene graphs for loop closure, or S-Graphs+ [7] [8], which couple in a situational graph modeled as a single optimizable factor graph, the 3D scene graph with the underlying SLAM graph. These works demonstrate that extracting semantic meaning from metric information and incorporating it with appropriate relational constraints such as walls, rooms, floors, etc., improve the overall precision of underlying SLAM while robustifying loop closures and generating meaningful environmental maps.
Recently, Hydra-Multi [9] and CURB-SG [10] took advantage of 3D scene graphs for collaborative SLAM in a centralized way. These approaches harness semantic-relational knowledge to exchange higher-level information between the agents to fortify inter-robot loop closures. However, these methods, apart from being centralized, explicitly add loop closure constraints to underlying robot poses instead of implicitly constraining the same semantic entities detected by different agents.
Thus in this work, we present Multi S-Graphs, a LiDAR-based distributed CSLAM system for structured environments harnessing the potential of the multi-layered Situational Graphs (S-Graphs), a single optimizable factor graph that combines a pose graph (i.e. with keyframes) and a 3D scene graph with semantic-relational concepts, to generate precise environmental maps and locate the different robots in this map while exchanging minimum information between robots. To the best of our knowledge, we present the first distributed collaborative SLAM system based on optimizable situational graphs. Fig. 1 shows the Multi S-Graphs executed on a real experiment.
@@
    A hybrid room-based descriptor combining raw point cloud information with semantic and hierarchical knowledge to identify inter-robot loop closures.
@@ II-A Multi-Agent SLAM
Figure 2: Multi S-Graph architecture viewed from $Robot_{i}$ perspective. Odometry measures and 3D Lidar data are the inputs that the robotic platform provides. Each robot executes its own Local S-Graph system and generates the Room Descriptors and the Distilled S-Graphs that are exchanged with the rest of the robots. Messages exchanged between robots are represented in gray. From the point of view of the $Robot_{i}$ the rest of the agents provide the Distilled graphs and the Room descriptors that are used to generate the Collaborative S-Graph. Information provided by the rest of the robots appears in green color.
@@
II-B Semantic Information in CSLAM
One of the challenges in CSLAM is combining the local maps of each robot into one global one when the initial pose of each one is unknown. Single-robot SLAM approaches like [24],[25],[26], utilize semantic object detection, enhancing the robustness and accuracy by reducing the loop closure search space. Motivated by this, some CSLAM approaches like [27],[28] utilize object-level semantics to robustify inter-robot map merging. In terms of map fusion, [29] presented a framework for efficient map-merging in semantic object-based SLAM detecting duplicate constellations of landmarks. In [30], the authors utilized semantic place recognition on sparse representations of tree trunks, represented by cylinder models extracted from a LIDAR sensor, to calculate relative transformations between multiple robots. All these works do not utilize hierarchical information along with semantic information to enhance inter-robot map merging as well as require densely populated indoor scenes with relevant objects.
II-C Multi-Agent 3D Scene Graphs
@@ III System Overview
1. Local S-Graph Generation (Section IV). To generate the local S-Graph for each robot, we use S-Graphs+ [8] which runs onboard each one using the odometry and 3D LiDAR measurements of the robot. It also estimates the position and trajectory followed by the robot while generating it. Like S-Graphs+, our system can work with any external odometry source.
@@
3. Room-Based Inter-Robot Loop Closure (Section VI). Relying only on the generated semantic graph for loop closure can generate alignment errors in the presence of symmetry in the environment. In order to combine the raw information of the point clouds and take advantage of the high-level semantic information that each room contains, we generate a hybrid descriptor for each room, a Room Descriptor. This module is in charge of two tasks: generating the descriptors and finding appropriate room matches.
@@ IV Local Situational Graph
We will use the four-wall rooms for the S-Graph distillation (Section V) and the Room-based loop closure (Section VI) because the boundaries and dimensions of these rooms can be easily extracted from the walls that constitute them.
@@ VI-A Room descriptor generator
To generate these descriptors, we rely on the Scan Context descriptor [21]. This is an egocentric, yaw-invariant, point cloud descriptor that has demonstrated state-of-the-art results, given its simplicity and fast convergence time. The main drawback of this descriptor is that it is sensitive to linear motions. To address this issue, we propose to generate Scan Contexts based on a Room Centric point cloud, which reduces the sensitivity to linear motion and strengthens descriptor matching.
In this work, we exploit the semantic and hierarchical information from Local S-Graph to obtain the room center and its boundaries to extract all the keyframes obtained from within the room and generate its corresponding Room Centric point cloud.
Each room keyframe $R_{k_{i}}$ can be expressed as:
@@ VI-B Loop Closure detection
Fine Alignment: Whenever a match is found between the robot and other agents’ room, it tries to obtain an improved transform by asking for the Room Centric point cloud that originates the Scan Context and using a VGICP[31] registration algorithm.
If a match is found, the room match with the relative transformation between rooms is sent into the following module (Section VII) to connect the different graphs.
Scan Context and ICP match validity depend on the $SC_{th}$ and $ICP_{th}$ thresholds respectively. The first one regulates the volume of point clouds sent over the network, while the second adjusts the match precision. It is crucial for Room Descriptors to remain unique and constant over time for accurate matches, as identical rooms or changing environments can lead to erroneous matches.
@@ VII Collaborative Situational Graph
Collaborative Graph: If a room match is found using the room matching step (Section VI), we formulate the factors between the local s-graph and the distilled s-graphs in the 2 steps. In the first step, a room-to-room factor is created for the matched rooms. For the given matched rooms, lower-level wall-to-wall factors are established, which robustify the alignment of the graphs. As seen in Fig 2, graphs are merged at rooms, walls, and floor levels. However, since floor nodes connect all the rooms at the same height, no explicit floor-to-floor factors are required between the floor nodes of the robots. Whenever local S-Graphs and external distilled S-Graphs are connected with appropriate factors, a common optimization is performed similarly to [8], leading into an optimized Collaborative S-Graph and, consequently, an optimized state estimation of the robot. After this optimization, the external graphs are aligned with the robot’s local coordinate frame. The new information provided by the other robots can be integrated into the collaborative graph, using the same coordinated frame as the local robot for representing the different entities of the external distilled graphs, as can be seen in Fig. 1. Subsequently, if a new room match is found, the same two-step room factors are formulated, increasing the number of factors between graphs and refining the optimized graph.
The Local S-Graph generated by S-Graphs+ is being updated continuously, adding new nodes into the graph and including intra-robot loop closures at the four hierarchical levels as in [7]. As the Collaborative Graph includes this local graph, these updates are also integrated into it.
@@ VIII Experimental Validation
VIII-A Methodology
(a) C1F0s
(b) C1F2s
(c) SE1
Figure 4: Simulated experiments, showing the trajectory and the point cloud map generated for each robot.
Real-world Data. The datasets utilized in our experiments consist of a collection of in-house datasets acquired from a construction site: C1F1, C1F2, C3F1, C3F2, and C2F2. In addition, an experiment with 3 robots was performed on the C2F2 stage enabled by the size of the scenario. For these experiments we relied on the robot encoders to generate the odometry data. Unfortunately, our real-world datasets do not include IMU data and consequently, algorithms that rely heavily on IMU, such as DCL-SLAM and Disco-SLAM, could not be evaluated.
(a) C1F1
(b) C1F2
(c) C3F1
(d) C3F2
(e) C2F2
VIII-B Experimental Setup
The robot used for generating the datasets is a Boston Dynamics Spot with a Velodyne VLP-16 onboard. The experiments were carried out using a laptop computer with an Intel i7-10870H (8 cores, 2.2 GHz) with 32 GB of RAM memory. Multi S-Graphs is developed under S-Graphs+[8] using ROS 2 for implementing the algorithms. The room matching parameters have been tuned to give a robust performance in the experimental environments, obtaining values of $SC_{th}=0.35$ and $ICP_{th}=0.07$.
VIII-C Results and Discussions
Simulated Data.
Fig. 4 shows the qualitative results of Multi S-Graphs obtained during the simulated experiments. The outcomes presented in Table I show that Multi S-Graph algorithm exhibits a lower Average Trajectory Error (ATE) for the aggregated robot trajectory compared with the selected state-of-the-art (SOTA) CSLAM algorithms. Our approach outperformed all other decentralized algorithms, reducing the accuracy error by more than $90\%$ ($93.7\%$, $94.5\%$, and $98.4\%$ for DCL-SLAM, Disco-SLAM, and Swarm-SLAM respectively). It also achieved better accuracy than LAMP 2.0, with $72.9\%$ less error, which requires precise initial poses of the robots. The lack of features in the simulation environment causes measurement errors to adversely affect optimization. The extraction of semantic information specific to this type of environment by our algorithm overcomes this source of error.
@@ VIII-C Results and Discussions
Ablation Study. The results of this study are collected in Table V. We can see that the utilization of hierarchical optimization, which establishes relationships between the walls of two rooms among agents (Rooms and Planes), means a significant improvement of accuracy, reducing the ATE by $25.6\%$ with an increase in the amount of data by $77\%$. On the other hand, the use of Fine alignment improves the performance by $73.6\%$ and $25.6\%$ over the Rooms and Rooms and Walls configurations, respectively, with a 40 times higher data exchange requirement. In general, fine alignment techniques perform better compared to their non-fine alignment counterparts, primarily due to the limited angular resolution of the Scan Context, which is insufficient for precise alignment (see Section VI). Furthermore, the use of wall-to-wall factors contributes to alignment improvement, consequently significantly enhancing overall performance. Consequently, the best performance is obtained by combining the semantic information extracted from the walls with fine alignment using the Room Centric point clouds, Rooms and Walls with Fine alignment, improving the performance by up to $74.7\%$. However, due to the larger amount of data exchanged required, when communication channels are limited, Rooms and Walls, which only use semantic information and avoid sending point cloud messages, may be an interesting solution without a significant decrease in accuracy.
@@
Figure 6: Accumulated data shared between agents, consisting of Walls Nodes (Wall GN) and Room Nodes (Room GN) for the Semantic Graph, and the Room Scan Contexts (Room SC) and Room Centric point clouds (Room PC).
@@ IX Conclusions and Future Work
In this work a distributed CSLAM system is presented, leveraging the semantic-relational features extracted by the S-Graphs [7], [8], in order to filter and reduce the amount of data that has to be transmitted between robots while outperforming state-of-the-art mapping and positioning accuracy in structured environments. We proposed a novel Room Descriptor that fuses the room structural elements with the fine-grained point cloud information, reducing the number of candidates for loop closure, without requiring the identification of objects present in the scene or exchanging keyframes between agents.
We tested our algorithm for a map generation task in multiple scenarios, simulated and real, outperforming other state-of-the-art lidar-based CSLAM algorithms in terms of precision and bandwidth, reducing the trajectory and map reconstruction error by more than $90\%$ while saving $95\%$ of the data exchanged between robots.
The main drawback is the dependence on semantic-relational structures, defined in our case as four-walled rooms in building-like environments, which are not always easy to extract. The room matching relies on these room structures and the detection of different static features in them, so having a non-feature environment or changing environment can cause the room matching not to work properly.
For future work, a more generic way of detecting different types of rooms and finding alignment between Situational Graphs should be found to increase the variety of scenarios in which Multi S-Graphs can be used. Also, a validation of the system running onboard the robots must be performed to test the real-time capabilities of the system.